The following data has been provided by your manufacturing department to estimate the standard cost of producing a particular product. Direct Labor - $20 per hour. The first unit took 0.6 hours of direct labor but you want to base the cost estimate on the number of hours needed for the 25th unit. You expect a 75% learning curve to apply to the direct labor hours. Overhead-50% of Direct Labor Material $43.75 per 25 units What is your estimate of the cost to produce each unit? If your company wishes to make 20% of the manufacturing cost as profit, what should the product be sold for?
Answer #1

Solution: Learning curve equation:

Y = aX ^b

Y = 0.6 × 25^(-0.415)   [Since b = In(0.75) / In(2) = - 0.415]

Y = 0.15776

Y = 0.16 (by rounding)

Standard cost per unit

Particulars

Amount $

Material ($43.75 / 25 =)

1.75

Direct labor ($20 × 0.16 =)

3.20

Overhead ($3.20 × 50% =)

1.60

Cost of producing each unit

6.55

+ Profit margin ($6.55 × 20% =)

1.31

Sale price

7.86

Answer: Standard cost is $6.55; Sale price is $7.86